The Gibsons invested $7000, part at 6% and the rest at 10% simple annual intrest. If they had reversed the amounts invested at each rate, they would have earned $120 less in annual intrest. How much was invested at each rate?

Answer by stanbon(75887)

The Gibsons invested $7000, part at 6% and the rest at 10% simple annual intrest. If they had reversed the amounts invested at each rate, they would have earned $120 less in annual interest. How much was invested at each rate?
----------------------------------
Original Int: 0.06x + 0.10(7000-x)
---
Reversed Int: 0.10x + 0.06(7000-x)
---
Equation:
original - reversed = 120
-----
[0.06x + 0.10(7000-x)] - [0.10x + 0.06(7000-x)] = 120
---
[-0.04x + 0.04(7000-x)] = 120
---
-0.04x + 0.04*7000 - 0.04x = 120
----
-0.08x = 120-0.04*7000
---
-0.08x = -160
---
x = $2000 (amt. invested at 6%)
---
7000-x = $5000 (amt. invested at 10%)
=========================================
Cheers,
Stan H.
===========